# A Critical Assessment of Visual Sound Source Localization Models Including Negative Audio

### v2

@@ Abstract
Although substantial progress has been made in VSSL in the last years regarding audio mixtures [7, 11, 14, 17, 19], only a few attempted to work with silent objects [8, 13] or offscreen sounds [13]. Most importantly, the majority of existing works evaluate the models with just positive audio examples, that is, with sounds (in a mixture or not) produced by objects that are visible in the image. Except for [16], no analysis is performed using negative audio, i.e. audio sources that are not visible. By evaluating with only positive cases, current approaches fail to represent real-world audio-visual scenarios and risk drawing misleading conclusions about how models interpret the audio input. While [16] considers silence and offscreen sounds, it does not provide metrics to analyse the negative sounds separately from the positive ones making it difficult to understand the model’s behaviour, and the proportion of audio-visual pairs with a negative audio is small compared to those with positive audio. Another shortcoming of current methods is that when evaluating, they either assume that the object is visible [21] and/or that their size is known [22], which is typically not the case in practice.
@@ II-3 F and F
With the aim of automatically choosing an adequate universal threshold for each model, we analyse in Fig. 1 the distribution of the maximum value in the audio-visual similarity map (i.e. before binarization), both for the samples in VGG-SS and IS3 test sets. Four types of audio samples for each image in the test set are used: positive, silence, noise and offscreen. We focus only on the maximum value because, in a positive case, it should exceed the threshold, while in a negative case, it should not (since no pixel in the image “sounds”). Fig. 1 shows a boxplot of the distribution of the maximum value, where the colored area inside the box represents the Inter Quartile Range (IQR), i.e. the second and third quartiles, where 50% of the values are concentrated. This figure illustrates the discriminatory power of each model between positive and negative audio. We analyze the following VSSL models: RCGrad [24], LVS [3], EZ-VSL [15], SLAVC [16], SSL-TIE [12] and SSL-Align [22].
As it can be observed in Fig. 1, most models present a strong overlap in the range of maximum similarities for the cases of positive, silence and offscreen audio, thus making it difficult to choose a universal threshold in which the negative audios will not be activated in the localization map. SSL-TIE and SSL-Align are the ones that present a higher discriminatory power. SSL-TIE is the most discriminative model in VGG-SS. In IS3, SSL-TIE is good at discriminating silence and noise but less good in case of offscreen. On the other hand, SSL-Align is good at discriminating noise and offscreen in IS3 but less good at silence. In light of the results in Fig. 1, we decide to set the universal localization threshold for each model as the highest 75th percentile, or 3rd quartile, of the maximum similarity values of the three types of negative audio. This value is computed in the extended test set of VGG-SS and is shown in Fig. 1 (both in numbers and as a horizontal line). We use this threshold to binarize the audio-visual similarity maps and thus obtain the localization map used in all metrics. This threshold used in the extended test sets of both VGG-SS and IS3; in this latter case with the aim of assessing if it generalizes. The threshold is computed differently for RCGrad model, as it uses gradCAM, with output values normalized in the interval . In this case, we use a threshold of .
@@ III Experiments
We conducted our experiments on the extended test sets of VGG-SS [3] and IS3 [22] datasets. Table I contains the results on the positive and each of the three negative audio inputs as well as the global metrics for both types of inputs, the best results are in bold and the second best underlined. All metrics except for two have been computed with our proposed universal threshold, U. th., (introduced in Section II). On the other hand, cIoU Adap. and AUC Adap. use the adaptive threshold proposed in [22]. In the case of positive audio input and IS3 test set, the first and second best models are SSL-Align and SSL-TIE, respectively, regardless of the threshold used to get the localization maps. In contrast, in VGG-SS, the best model with the universal threshold is SSL-TIE and SLAVC second. With the adaptive threshold, i.e. the one that uses the prior knowledge of the sounding object’s size, the best results are achieved by SSL-Align for cIoU and by SSL-TIE for AUC. EZ-VSL is the best in silence in VGG-SS and second best in IS3, while SSL-TIE is best in IS3 and second best in VGG-SS. LVS is best in noise and offscreen for both datasets. SSL-Align is second best in noise in VGG-SS and SSL-TIE in IS3. EZ-VSL is second best in offscreen in IS3 and SSL-Align in VGG-SS. Although, LVS and EZ-VSL have poor results in positive audio inputs, thus providing the worst global metrics. The global performance for both positive and negative audio is measured by F and F. We observe that best overall model in VGG-SS is SSL-TIE and in IS3 is SSL-Align, while SSL-TIE is the second best in IS3 and SLAVC in VGG-SS. The results in Table I are consistent with the analysis shown in Figure 1, where the best models according to the global metrics are the ones with less overlap in the IQR of maximum similarity values in the positive and negative cases , i.e. the more discriminatory ones.
A qualitative example is shown in Figure 2. Each row represents a different model, and each column corresponds to different audio inputs paired with the same image. The example is from IS3 and there are five different positive audio files: ‘man’, ‘ukulele’ and the three types of negative audio samples. To show the effect of the universal threshold we set to zero the audio-visual similarity values below that threshold. The resulting values are then normalized to the range and they are combined with the original image. Although SSL-Align is the best overall model in IS3, it fails in this example with the audio corresponding to ‘man’ and ‘silence’. SSL-TIE correctly localizes part of the regions corresponding to both positive audio and produces empty maps for the negative audio samples. EZ-VSL, SLAVC and FNAC produce empty maps for the negatives as well as for the man. In contrast, LVS highlights almost the same region for three different audio inputs: ‘man’, ‘ukulele’ and ‘silence’. Finally, RCGrad outputs the same localization regardless of the audio input.
@@ IV Conclusion
In this paper we assessed the performance of various SOTA Visual Sound Source Localization models, not only in the positive audio case, which is the common practice, but also complementing the evaluation on three types of negative audio inputs: silence, noise and offscreen sound. We proposed metrics to assess the localization for negative audio samples and two new metrics (F and F) which evaluate the performance globally in both positive and negative cases. Additionally, we have established a universal threshold for performing sound localization in an in-the-wild scenario, where no prior knowledge about the sounding object visibility and size is assumed. Through qualitative and quantitative testing across the VGG-SS and IS3 extended test sets, we identified the models that better localize the regions where the sound comes from, while being robust to negative audio inputs, although no model was best in all cases.

### v3

@@ Abstract
Although substantial progress has been made in VSSL in the last years regarding audio mixtures [7, 11, 14, 17, 19], only a few attempted to work with silent objects [8, 13] or offscreen sounds [13]. Most importantly, the majority of existing works evaluate the models with just positive audio examples, that is, with sounds (in a mixture or not) produced by objects that are visible in the image. Except for [16], no analysis is performed using negative audio, i.e. audio sources that are not visible. By evaluating with only positive cases, current approaches fail to represent real-world audio-visual scenarios and risk drawing misleading conclusions about how models interpret the audio input. While [16] considers silence and offscreen sounds, it does not provide metrics to analyse the negative sounds separately from the positive ones, making it difficult to understand the model’s behaviour, and the proportion of audio-visual pairs with a negative audio is small compared to those with positive audio. Another shortcoming of current methods is that when evaluating, they either assume that the object is visible [21] and/or that their size is known [22], which is typically not the case in practice.
@@ II-3 F and F
With the aim of automatically choosing an adequate universal threshold for each model, we analyse in Fig. 1 the distribution of the maximum value in the audio-visual similarity map (i.e. before binarization), both for the samples in VGG-SS and IS3 test sets. Four types of audio samples for each image in the test set are used: positive, silence, noise and offscreen. We focus only on the maximum value because, in a positive case, it should exceed the threshold, while in a negative case, it should not (since no pixel in the image “sounds”). Fig. 1 shows a boxplot of the distribution of the maximum value, where the colored area inside the box represents the Inter Quartile Range (IQR), i.e. the second and third quartiles, where 50% of the values are concentrated. This figure illustrates the discriminatory power of each model in distinguishing between positive and negative audio. We analyze the following VSSL models: RCGrad [24], LVS [3], EZ-VSL [15], SLAVC [16], SSL-TIE [12] and SSL-Align [22].
As it can be observed in Fig. 1, most models present a strong overlap in the range of maximum similarities for the cases of positive, silence and offscreen audio, thus making it difficult to choose a universal threshold in which the negative audios will not be activated in the localization map. SSL-TIE and SSL-Align are the ones that present a higher discriminatory power. SSL-TIE is the most discriminative model in VGG-SS. In IS3, SSL-TIE is good at discriminating silence and noise but less good in case of offscreen. On the other hand, SSL-Align is good at discriminating noise and offscreen in IS3 but less good at silence. In light of the results in Fig. 1, we decide to set the universal localization threshold for each model as the highest 75th percentile, or 3rd quartile, of the maximum similarity values of the three types of negative audio. This value is computed in the extended test set of VGG-SS and is shown in Fig. 1 (both in numbers and as a horizontal line). We use this threshold to binarize the audio-visual similarity maps and thus obtain the localization map used in all metrics. This threshold is used in the extended test sets of both VGG-SS and IS3; in this latter case with the aim of assessing if it generalizes. The threshold is computed differently for RCGrad model, as it uses gradCAM, with output values normalized in the interval . In this case, we use a threshold of .
@@ III Experiments
We conducted our experiments on the extended test sets of VGG-SS [3] and IS3 [22] datasets (for VGG-SS we use 3,983 out of 5,000 videos that were available on YouTube). Table I contains the results on the positive and each of the three negative audio inputs as well as the global metrics for both types of inputs, the best results are in bold and the second best underlined. All metrics except for two have been computed with our proposed universal threshold, U. th., (introduced in Section II). On the other hand, cIoU Adap. and AUC Adap. use the adaptive threshold proposed in [22]. In the case of positive audio input and IS3 test set, the first and second best models are SSL-Align and SSL-TIE, respectively, regardless of the threshold used to get the localization maps. In contrast, in VGG-SS, the best model with the universal threshold is SSL-TIE and SLAVC second. With the adaptive threshold, i.e. the one that uses the prior knowledge of the sounding object’s size, the best results are achieved by SSL-Align for cIoU and by SSL-TIE for AUC. EZ-VSL is the best in silence in VGG-SS and second best in IS3, while SSL-TIE is best in IS3 and second best in VGG-SS. LVS is best in noise and offscreen for both datasets. SSL-Align is second best in noise in VGG-SS and SSL-TIE in IS3. EZ-VSL is second best in offscreen in IS3 and SSL-Align in VGG-SS. Although, LVS and EZ-VSL have poor results in positive audio inputs, thus providing the worst global metrics. The global performance for both positive and negative audio is measured by F and F. We observe that best overall model in VGG-SS is SSL-TIE and in IS3 is SSL-Align, while SSL-TIE is the second best in IS3 and SLAVC in VGG-SS. The results in Table I are consistent with the analysis shown in Figure 1, where the best models according to the global metrics are the ones with less overlap in the IQR of maximum similarity values in the positive and negative cases , i.e. the more discriminatory ones.
A qualitative example is shown in Figure 2. Each row represents a different model, and each column corresponds to different audio inputs paired with the same image. The example is from IS3 and there are two different positive audio files: ‘man’, ‘ukulele’ and the three types of negative audio samples. To show the effect of the universal threshold we set to zero the audio-visual similarity values below that threshold. The resulting values are then normalized to the range and they are combined with the original image. Although SSL-Align is the best overall model in IS3, it fails in this example with the audio corresponding to ‘man’ and ‘silence’. SSL-TIE correctly localizes part of the regions corresponding to both positive audio and produces empty maps for the negative audio samples. EZ-VSL, SLAVC and FNAC produce empty maps for the negatives as well as for the man. In contrast, LVS highlights almost the same region for three different audio inputs: ‘man’, ‘ukulele’ and ‘silence’. Finally, RCGrad outputs the same localization regardless of the audio input.
@@ IV Conclusion
In this paper we assessed the performance of various SOTA Visual Sound Source Localization models, not only in the positive audio case, which is the common practice, but also complementing the evaluation on three types of negative audio inputs: silence, noise and offscreen sound. We proposed new metrics to assess the localization for negative audio samples and two new metrics (F and F) which evaluate the performance globally in both positive and negative cases. Additionally, we have established a universal threshold for performing sound localization in an in-the-wild scenario, where no prior knowledge about the sounding object visibility and size is assumed. Through qualitative and quantitative testing across the VGG-SS and IS3 extended test sets, we identified the models that better localize the regions where the sound comes from, while being robust to negative audio inputs, although no model was best in all cases.